If you have an aqueous solution that is 14.4 % Na3PO4 by mass, what is the molality of Na3PO4 in the solution?

Enter your answer in units of molality to three significant figures.
Answer #1

% m/m = (mass of solute / mass of solution) x 100

assume 100 g solution

14.4 % m/m = (W / 100) x 100

W / 100 = 0.144

W = 14.4 g

mass of solvent = mass of solution - mass of solute = 100 - 14.4 = 85.6 g

molality = (W / MW) (1000 / mass of solvent)

molality = (14.4 / 163.94) (1000 / 85.6)

molality = 1.03 m